OpenStudy (anonymous):

How do you convert from one unit to the other?? Liters to cm, Diameters to cm???? I have no clue where to start :/

OpenStudy (anonymous):

Usually you need a conversion rate, which will have have a unit on top and a unit on bottom. For example: \[ \frac{1000 mL}{1 L}\] (Since there are 1000 mL per L) You want the unit to which you are converting to be on top, and the unit from which you are converting to be on bottom. Remember that: \[ 1 mL = 1cm^3 \] So: \[ x L \cdot \frac{1000 mL}{1L} \cdot \frac{1 cm^3}{1 mL} = 1000\cdot x \ cm^3 \] Diameter is not a unit, but decameter and decimeter are.